z/20+1/12=z/15-1/10
Answer by tommyt3rd(5050)

z/20+1/12=z/15-1/10

z/20-z/20+1/12+1/10=z/15-z/20-1/10+1/10
5/60+6/60 = 4z/60-3z/60
11/60 = z/60

z=11